OpenStudy (anonymous):

Find the derivative of the given function:

OpenStudy (anonymous):

\[y=\frac{ (2x)^{2} }{ e ^{2x} }\]

OpenStudy (ipwnbunnies):

Use quotient rule.

OpenStudy (anonymous):

quotient property or quotient rule?

OpenStudy (ipwnbunnies):

Err, Idr. The quotient rule for derivatives: \[\frac{dy}{dx}[\frac{f(x)}{g(x)}] = \frac{g(x)*f'(x) - f(x)*g'(x)}{(g(x))^2}\]

OpenStudy (anonymous):

ehh, i'll let u know. thanks

OpenStudy (ipwnbunnies):

The bottom times the derivative of the top minus the top times the derivative of bottom. All over the bottom squared. It's catchy.

OpenStudy (anonymous):

its going to be hard with e^2x

OpenStudy (anonymous):

I am currently on logarithms

OpenStudy (anonymous):

I remember this formula; in the beggining of the year

OpenStudy (anonymous):

I have a feeling it doesn't apply. :|

OpenStudy (ipwnbunnies):

\[\frac{d}{dx} [e^{u}] = e^{u}*\frac{du}{dx}\]

OpenStudy (ipwnbunnies):

Chain rule on exponential function with 'e' is simple.

OpenStudy (anonymous):

@iPwnBunnies Help me with a hint?

OpenStudy (anonymous):

I've never performed this before

OpenStudy (ipwnbunnies):

Let's find the first term in the quotient rule: g(x)*f'(x)

OpenStudy (anonymous):

(e^2x)(4x)

OpenStudy (ipwnbunnies):

Try again. The numerator simplifies to 4x^2.

OpenStudy (anonymous):

the derivative of 2x^2 simpliesfied to 4x

OpenStudy (ipwnbunnies):

(2x)^2 = 4x^2

OpenStudy (anonymous):

okay. i got that

OpenStudy (ipwnbunnies):

What's the first term now? lol

OpenStudy (anonymous):

(e^2x)(4x^2)

OpenStudy (ipwnbunnies):

No. That's what the numerator in the original function simplifies to.

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

8x^3e+8xe^2x

OpenStudy (anonymous):

did you expect me to find the derivative of e^2x

OpenStudy (ipwnbunnies):

Whoa, let's go back to our simplifed original function. \[y = \frac{4x^{2}}{e^{2x}}\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

kinda lost

OpenStudy (ipwnbunnies):

Ahh crap. Didn't realize this was gonna take this long. I'll show you the derivative of e^2x.

OpenStudy (anonymous):

e=10?

OpenStudy (ipwnbunnies):

In e^(2x), let u = 2x, and du = 2 dx \[\frac{d}{dx}[e^{u}] = e^{u}*\frac{du}{dx}\]

OpenStudy (ipwnbunnies):

As you can see, you keep the e^(2x), then multiply it by the derivative of its power. The derivative of e^(2x) = e^(2x) * 2 = 2e^(2x)

OpenStudy (anonymous):

what about dx?

OpenStudy (ipwnbunnies):

It's in the du/dx

OpenStudy (ipwnbunnies):

du = 2 dx, du/dx = 2

OpenStudy (anonymous):

okay... i see

OpenStudy (ipwnbunnies):

Now, if you follow the quotient rule for derivatives, you should be good. Sorry man. G2g.

OpenStudy (anonymous):

thanks a lot